High-Resolution Camera Report from Liam Pedersen

Condensation Problem:

Moisture in heated camera enclosure on Pan Tilt mounted colour camera condenses and freezes on front glass pane:

Probable cause is snow entering through small holes in camera enclosure, melting due to heaters, and condensing and freezing on cold front plate. Holes are very small, largest approx 1 cm2.

Duct tape (lots of it) and rice. Use duct-tape to seal all openings, remove all interior snow, and put in bags of rice to absorb water vapour in enclosure. (No other dessicant was available).

Has worked so far. Should mention that convenient camera access is need in future. This enclosure was almost TOO robust.

Camera Scaling Calibration Experiment, Perfromed 11/10/98

Angular field of view (fov) of camera pixels various zooms had to be deduced experimentally by imaging a calibration target of known dimensions at various distances from the camera and zoom settings. This information is combined with the distance from the camera to the image subject to compute the scaling (mm/pixel) of the acquired image and thus compute the dimensions of the subject.

Complications occur because the camera zoom is measured by an encoder that outputs numbers whos relationship to the actual zoom value (in mm) is unknown. Furthermore, distance to object not known exactly, but is deduced from direction camera is pointing, assuming that the robot is on flat ground.

Missing Colour Channels

The camera occasionally only acquires 2 of the 3 colour channels (R,G,B):

Currently, the blue channel is flaky (see figure above), typically 20 times less bright than the other channels. Suspicions are that the Mutech frame grabber is at fault, since cables have been replaced, different cameras tried. Furthermore, the corrupted channel is not consistent. Sometimes the red or green channels are corrupted instead.

The rock classification code may be modified to deal with the loss of the blue channel, at the expense of accuracy. It cannot be changed to cope with the loss of the other channels. Intentions are to use Kodak handheld digital camera to get supplementary images of all rocks for post validation of classification code.

Expedition Report from William Cassidy

26. Nov. 21, 1998, Saturday: Patriot Hills Camp

The high winds continued overnight, piling up 5-foot high drifts in camp. Still windy this morning and into the afternoon. The team is ready to go out to Nomad as soon as it becomes feasible but the very bad weather makes work outside impossible. We are holed up indoors, waiting for a change. Nothing else to report.